Solve each equation on the interval 0 1-cos theta= 1/2
Answer by lwsshak3(11628)

Solve each equation on the interval 0<2pi
1-cos theta= 1/2
cos theta=1-1/2=1/2
theta=π/3 and 5π/3 (reference angle π/3 in quadrants I and IV where cos>0)